OpenStudy (anonymous):

The Fussy Food Deli serves a simple fruit salad with strawberries and bananas. If strawberries cost $2.30/lb and bannas cost $.30/lb how many pound of strawberries and how many pound of bananas should be used to make a 5 pound blend that cost $1.70/lb

OpenStudy (anonymous):

You'll need to write formulas to express the word problem, then solve. You can write the following formulas... S = 1 lb of strawberries B = 1 lb of bananas S + B = 5 2.3*S + .3*B = 1.7*5 Then solve one of the formulas for S, and plug that value into the other formula and solve for B. That answer for B will tell you how many pounds of bananas to use.